OpenStudy (anonymous):

Solve 2 3/4 + x + 3 1/2 = 10

OpenStudy (catlover5925):

X = 3 and 3/4ths...if you need how I did it let me know

OpenStudy (anonymous):

\[ 2 + \frac {3}{4} + x + 3 + \frac {1}{2} = 10 \implies 5 + \frac {5}{4} + x = 10 \implies \frac {5}{4} + x = 5. \]Can you finish from here?

OpenStudy (anonymous):

thank you